hartnn (hartnn):

Half-range sine series. Check my work, Either I have done some mistake or its a tricky deduction! How would i deduce that ?

OpenStudy (anonymous):

I can't see the problem...

hartnn (hartnn):

the problem is i have a "n^3" in the denominator whereas the deduction requires 1/3^2 , 1/5^2 ....so 1/n^2

hartnn (hartnn):

@phi @ganeshie8

OpenStudy (phi):

my version of word can't read your file.

hartnn (hartnn):

OMG, did they just do a typing error and in the question there actually should be cubes and not squares ?? still i get this, 1/12 missing...

hartnn (hartnn):

\(π^3/32=1/1^3 -1/3^3 +1/5^3 -1/7^3 +..… \)

hartnn (hartnn):

i think they meant 1/1^2 which is still incorrect

hartnn (hartnn):

and \(π^3/32=1/1^3 -1/3^3 +1/5^3 -1/7^3 +..…\) is what they want

OpenStudy (phi):

Yes, your work is good, and their summation is something entirely different as you probably figured out.

OpenStudy (phi):

\[ \frac{\pi^3}{32} = \sum_{k=0}^\infty \frac{(-1)^k}{(2k+1)^3} \]

hartnn (hartnn):

thanks for the confirmation! :)